# Exploring strong electronic correlations in the breathing kagome metal Fe3Sn

### v1

@@ V Conclusions
In conclusion, we conducted a comprehensive theoretical investigation of the electronic structure and magnetic properties of the three-dimensional breathing-kagome ferromagnet Fe3Sn, using DFT, DFT+DMFT, and DFT+U. We calculated both intralayer and interlayer magnetic exchange interactions between the magnetic moments at the Fe sites, along with their respective band structures (spectral functions). Our results show that DFT+DMFT can be used to accurately reproduce the experimentally reported value of the MAE and to determine the most important contributions to the ferromagnetic coupling, including the presence of a 2-3% anisotropic contribution in the form of the Dzyaloshinskii-Moriya interaction. When the local coulomb interaction is considered explicitly, via DFT+DMFT and DFT+U, the quasiparticle spectra undergo significant changes, with a series of non-dispersive bands moving closer to the Fermi energy and even crossing it. Albeit this last conclusion depends on the particular choice of Coulomb interaction parameters, which is difficult to ascertain from first principles, these results point to a high tunability of the magnetic response, already reported for Fe3Sn2 Tanaka et al. (2020), and in line with the high sensitivity to doping and structural variations Vekilova et al. (2019); Fayyazi et al. (2019). Overall, topological Lifshitz transitions are expected to be tunable by small stimuli, such as strain and doping, which may affect the anomalous Hall and Nersnt effects substantially.
@@ Acknowledgments
We would like to thank H.S. Kim, I.E. Brumboiu, L. Prodan, and V. Tsurkan for insightful discussions and technical help. Computational work was performed on resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We also acknowledge Polish high-performance computing infrastructure PLGrid for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through PLL/2023/04/016450. I.D.M. also acknowledges financial support from the European Research Council (ERC), Synergy Grant FASTCORR, Project No. 854843. This research is also part of the project No. 2022/45/P/ST3/04247 co-funded by the National Science Center and the European Union’s Horizon 2020 research and innovation program under the Marie Skodowksa-Curie grant agreement no. 945339.

### v2

@@ V Conclusions
In conclusion, we conducted a comprehensive theoretical investigation of the electronic structure and magnetic properties of the three-dimensional breathing-kagome ferromagnet Fe3Sn, using DFT, DFT+DMFT, and DFT+U. We calculated both intralayer and interlayer magnetic exchange interactions between the magnetic moments at the Fe sites, along with their respective band structures (spectral functions). Our results show that DFT+DMFT can be used to accurately reproduce the experimentally reported value of the MAE and to determine the most important contributions to the ferromagnetic coupling, including the presence of a 2-3% anisotropic contribution in the form of the Dzyaloshinskii-Moriya interaction. When the local coulomb interaction is considered explicitly, via DFT+DMFT and DFT+U, the quasiparticle spectra undergo significant changes, with a series of non-dispersive bands moving closer to the Fermi energy and even crossing it. Albeit this last conclusion depends on the particular choice of Coulomb interaction parameters, which is difficult to ascertain from first principles, these results point to a high tunability of the magnetic response, already reported for Fe3Sn2 [18], and in line with the high sensitivity to doping and structural variations [25, 23]. Overall, topological Lifshitz transitions are expected to be tunable by small stimuli, such as strain and doping, which may affect the anomalous Hall and Nernst effects substantially. A key prerequisite for clearly observing this tunability is that the system does not undergo a significant phase transition, as e.g. due to emergent structural distortions or changes in the magnetic ground state.
@@ Acknowledgments
We would like to thank H.S. Kim, I.E. Brumboiu, L. Prodan, and V. Tsurkan for insightful discussions and technical help. Computational work was performed on resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We also acknowledge Polish high-performance computing infrastructure PLGrid for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through PLL/2023/04/016450. I.D.M. acknowledges financial support from the European Research Council (ERC), Synergy Grant FASTCORR, Project No. 854843, as well as from the STINT Mobility Grant for Internationalization (grant no. MG2022-9386). L.C. acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through TRR 360, project no. 492547816. This research is also part of the project No. 2022/45/P/ST3/04247 co-funded by the National Science Center of Poland and the European Union’s Horizon 2020 research and innovation program under the Marie Skodowksa-Curie grant agreement no. 945339.